### v1

# Variational Self-Supervised Learning

Abstract
We present Variational Self-Supervised Learning (VSSL), a novel framework that combines variational inference with self-supervised learning to enable efficient, decoder-free representation learning. Unlike traditional VAEs that rely on input reconstruction via a decoder, VSSL symmetrically couples two encoders with Gaussian outputs. A momentum-updated teacher network defines a dynamic, data-dependent prior, while the student encoder produces an approximate posterior from augmented views. The reconstruction term in the ELBO is replaced with a cross-view denoising objective, while preserving the analytical tractability of Gaussian KL divergence terms. We further introduce cosine-based formulations of KL and log-likelihood terms to enhance semantic alignment in high-dimensional latent spaces. Experiments on CIFAR-10, CIFAR-100, and ImageNet-100 show that VSSL achieves competitive or superior performance to leading self-supervised methods, including BYOL and MoCo V3. VSSL offers a scalable, probabilistically grounded approach to learning transferable representations without generative reconstruction, bridging the gap between variational modeling and modern self-supervised techniques.
Index Terms:
self-supervised learning, variational inference, representation learning, encoder-only models
I Introduction
Variational inference has become a foundational paradigm in machine learning, enabling scalable approximations of posterior distributions in latent variable models. The variational autoencoder (VAE), introduced by Kingma and Welling (2013), exemplifies this approach by optimizing the Evidence Lower Bound (ELBO), which balances a reconstruction term—computed via a decoder—with a Kullback-Leibler (KL) divergence regularizer aligning the approximate posterior to a prior [1]. While effective, the decoder’s role in reconstructing input data introduces computational complexity and assumes reconstruction as a prerequisite for meaningful representations, a constraint that may not always be necessary. In parallel, self-supervised learning has gained prominence for its ability to extract high-quality features without explicit reconstruction, often outperforming supervised methods in representation learning tasks. For instance, contrastive methods like SimCLR [2] leverage mutual information between augmented views of data to learn robust features, while predictive frameworks such as BYOL [3] achieve similar success by aligning representations across networks without negative samples. It is well-established that self-supervised methods excel at learning quality features, capturing intricate patterns and transferable representations more effectively than traditional supervised or generative approaches [4].
Motivated by these insights, we propose a novel variational self-supervised learning framework that eschews the decoder entirely, instead symmetrically coupling two encoders with Gaussian-distributed outputs. In our approach, each encoder generates an approximate posterior conditioned on the input and predicts the prior of the other’s latent representation, derived from a momentum-based network. This formulation retains the analytical tractability of Gaussian KL divergence [5], while replacing the reconstruction objective with a cross-predictive task. By integrating variational principles with self-supervised symmetry and momentum encoder, our method offers a decoder-free alternative that aligns with the strengths of recent representation learning paradigms.
[…]
Figure 1: Directed graphical model for the VSSL framework. Observations $x_{1}$ and $x_{2}$ are encoded via parameterized inference networks $\theta_{t}$ and $\theta_{s}$, producing the latent representation $z$. The student path $q_{\theta_{s}}(z\mid x_{2})$ updates the teacher path $q_{\theta_{t}}(z\mid x_{1})$ through exponential moving average (EMA). A denoising network $p_{\phi}(\hat{z}\mid z)$ refines the latent, producing a denoised representation $\hat{z}$ for self-supervised learning objectives.
II Related Work
Variational inference has emerged as a cornerstone of probabilistic modeling in machine learning, offering a scalable framework for approximating posterior distributions in latent variable models. The variational autoencoder (VAE), introduced by Kingma and Welling [1], serves as a seminal example, optimizing the Evidence Lower Bound (ELBO) by balancing a reconstruction term with a Kullback-Leibler (KL) divergence regularizer. The VAE’s reliance on a decoder to reconstruct input data, while effective for generative tasks, introduces computational overhead and assumes reconstruction is essential for learning meaningful representations—a constraint that may not always align with representation learning objectives. Subsequent advancements, such as $\beta$-VAE [6] and InfoVAE [7], have refined this framework by adjusting the trade-off between reconstruction fidelity and latent regularization, yet they largely retain the decoder-centric structure.
In parallel, self-supervised learning has gained traction as a powerful paradigm for learning high-quality representations without explicit supervision or reconstruction. Contrastive methods, like SimCLR [2], leverage mutual information maximization between augmented views of the same input to learn robust features, often outperforming supervised approaches in downstream tasks like image classification and transfer learning [4]. However, these methods typically require negative samples, which increases computational complexity. Predictive self-supervised frameworks, such as BYOL [3] and SimSiam [8], eliminate the need for negative samples by aligning representations across two networks—one often updated via momentum—demonstrating that reconstruction-free objectives can yield competitive representations. These approaches highlight the potential of encoder-only architectures, though they generally lack the probabilistic grounding provided by variational methods.
Efforts to bridge variational inference and self-supervised learning have been explored in prior work. For instance, Variational Predictive Coding [9] integrates predictive objectives into a variational framework, but it still typically assumes a generative component akin to a decoder. Similarly, Contrastive Variational Autoencoders [10] merge these concepts, yet often retain the decoder for reconstruction or related generative tasks. More recently, works like Barlow Twins [11] and VICReg [12] have proposed non-contrastive objectives based on redundancy reduction and variance preservation, aligning representations without reconstruction, though they do not explicitly adopt a variational formulation. Other self-supervised variational learning approaches [13, 14] combines contrastive learning with variational objectives. These methods underscore the shift toward decoder-free representation learning but often sacrifice the analytical tractability and probabilistic interpretability associated with variational inference.
Figure 2: Overview of the variational self-supervised learning framework for unsupervised representation learning using variational objectives. An original image is augmented into two views, $t_{i}$ and $t_{j}$, processed by a momentum-updated teacher network and a student network, respectively. Both networks encode the image into feature vectors and variational distributions parameterized by $\mu$ and $\log\sigma$. The teacher outputs serve as a prior for the student’s posterior via a KL divergence minimization. Gaussian sampling from the student posterior allows further processing through an autoencoder, enforcing consistency and regularization in the learned latent space.
Our proposed framework builds on these insights by fully eschewing the decoder and symmetrically coupling two encoders with Gaussian-distributed outputs, inspired by the momentum-based priors of BYOL [3]. Unlike traditional VAEs, we replace the reconstruction term with a cross-predictive task where each encoder’s approximate posterior predicts the prior of the other, leveraging the closed-form KL divergence for Gaussian distributions [5]. This approach aims to marry the probabilistic rigor of variational inference with the feature-learning strengths of self-supervised methods, distinguishing it from prior hybrid models that retain generative components or lack a symmetric encoder design. By eliminating the decoder while preserving analytical tractability, our method offers a potentially novel and efficient alternative for latent representation learning.
III Methodology
In this section, we detail the methodology for a novel variational self-supervised learning (VSSL) framework that eliminates the need for a decoder. Instead, it symmetrically couples two encoders with Gaussian-distributed outputs to learn meaningful latent representations. By integrating variational inference with self-supervised learning principles, this approach replaces the traditional reconstruction objective with a cross-predictive task, leveraging the analytical tractability of Gaussian distributions.
III-A Variational Formulation
The variational autoencoder (VAE) is a generative model designed to learn a probabilistic latent representation of data. For a given data point $x$, the VAE optimizes the Evidence Lower Bound (ELBO) on the marginal log-likelihood $\log p(x)$:
$\mathcal{L}(x)=\mathbb{E}_{q(z\mid x)}[\log p(x\mid z)]-D_{\mathrm{KL}}(q(z\mid x)\|p(z))$
Here:
  • $q(z\mid x)$: approximate posterior distribution, parameterized by an encoder network.
  • $p(x\mid z)$: likelihood, parameterized by a decoder network.
  • $p(z)$: prior over the latent variable $z$, typically a standard normal $\mathcal{N}(0,I)$.
The ELBO consists of two terms:
  1. 1.
    Reconstruction Term: $\mathbb{E}_{q(z\mid x)}[\log p(x\mid z)]$, encouraging accurate reconstruction of $x$ from $z$.
  2. 2.
    KL Divergence Term: $D_{\mathrm{KL}}(q(z\mid x)\|p(z))$, regularizing the posterior to align with the prior.
III-B Self-Supervised Learning Context
Self-supervised learning leverages supervisory signals from the data itself, often through multiple views or augmentations. In this work, we adapt the VAE framework to self-supervised learning using a teacher-student mechanism and a denoising task. Two views, $x_{1}$ and $x_{2}$ (e.g., augmentations of the same sample), are used to improve the robustness of the latent representations.
III-C EMA-Updated Teacher Network as Prior
In standard VAEs, the prior $p(z)$ is fixed as $\mathcal{N}(0,1)$. We introduce a dynamic, data-dependent prior using a teacher network with parameters $\theta_{t}$, updated via an exponential moving average (EMA) of the student encoder parameters $\theta_{s}$:
$\theta_{t}\leftarrow\tau\theta_{t}+(1-\tau)\theta_{s}$
where $\tau\in[0,1)$ is a decay factor. The teacher processes view $x_{1}$ to define the prior:
$p_{\theta_{t}}(z\mid x_{1})=\mathcal{N}(z\mid\mu_{t}(x_{1}),\sigma_{t}^{2}(x_{1}))$
The KL divergence in the ELBO becomes:
$D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
where $q_{\theta_{s}}(z\mid x_{2})=\mathcal{N}(z\mid\mu_{s}(x_{2}),\sigma_{s}^{2}(x_{2}))$ is the student posterior. This term encourages consistency between the student’s encoding of $x_{2}$ and the teacher’s prior from $x_{1}$, stabilized by the EMA update.
III-D Reconstruction Term with Denoiser Networks
We replace the standard reconstruction term with a denoising objective. Denoiser networks, parameterized by $\phi$, predict distributional parameters from view $x_{1}$:
$p_{\phi}(z\mid\mu_{1},\sigma_{1})=\mathcal{N}(z\mid\mu_{\phi}(x_{1}),\sigma_{\phi}^{2}(x_{1}))$
Letting $\mu_{1}=\mu_{\phi}(x_{1})$, $\sigma_{1}=\sigma_{\phi}(x_{1})$, the new reconstruction objective becomes:
$\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
Here, $z$ is sampled using the reparameterization trick: $z=\mu+\sigma\cdot\epsilon$, with $\epsilon\sim\mathcal{N}(0,1)$.
This jointly trains:
  • The student encoder, $q_{\theta_{s}}(z\mid x_{2})$, to map $x_{2}$ to latent space.
  • The denoiser networks, $p_{\phi}(z\mid\mu_{1},\sigma_{1})$, to infer latent likelihood from $x_{1}$.
This setup enforces probabilistic alignment between the latent codes of $x_{1}$ and $x_{2}$, avoiding direct input reconstruction.
III-E Derivation of the Self-Supervised ELBO
We derive the modified ELBO for this self-supervised framework. Given views $x_{1}$ and $x_{2}$, we maximize the likelihood $\log p_{\phi}(z\mid\mu_{1},\sigma_{1})$. Applying Jensen’s inequality gives:
$\log p_{\phi}(z\mid\mu_{1},\sigma_{1})$ $\geq\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad-D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
Thus, the modified ELBO becomes:
$\mathcal{L}(x_{1},x_{2})$ $=\mathbb{E}_{q_{\theta_{s}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad+\mathbb{E}_{q_{\theta_{t}}(z\mid x_{2})}[\log p_{\phi}(z\mid\mu_{1},\sigma_{1})]$
$\quad-D_{\mathrm{KL}}(q_{\theta_{s}}(z\mid x_{2})\|p_{\theta_{t}}(z\mid x_{1}))$
Likelihood Terms: Encourage consistency between $x_{2}$’s latent samples and $x_{1}$’s predicted distribution. KL Term: Regularizes the student’s posterior to align with the EMA-stabilized teacher prior.
III-F Implementation Details
Augmentation. VSSL  [^1] GitHub Repository: github.com/convergedmachine/vssl-solo-learn employs SimCLR-style augmentations [2], including random cropping to 224×224, horizontal flips, color jitter (brightness, contrast, saturation, hue), and optional grayscale conversion to enhance view diversity.
Architecture. The framework includes an online student network and a momentum-updated teacher network. Projectors, Predictors, and Denoisers: Each component consists of a two-layer MLP with batch normalization and ReLU activations, projecting input features into mean and variance parameters. A primary predictor further refines these representations, while two denoising networks independently predict mean and variance for the latent reconstruction objective.
Variational Modeling. Latent variables are modeled as Gaussian distributions with parameters constrained via the Softplus activation on variance terms. Stability is enhanced by adding noise proportional to $0.01\cdot\epsilon^{2}$, where $\epsilon\sim\mathcal{N}(0,1)$.
Training. Each training step involves:
  1. 1.
    Computing the student posterior and teacher prior.
  2. 2.
    Estimating the denoising likelihood from the latent features.
  3. 3.
    Calculating the KL divergence between the student and teacher distributions.
Gradients are used to update student parameters, while teacher parameters are updated using exponential moving average (EMA) of the student weights.
Cosine-Based KL and Log-Likelihood. To enhance semantic alignment in high-dimensional latent spaces, we propose cosine-based alternatives to classical KL divergence and log-likelihood formulations.
Cosine-Based KL Divergence:
$\tilde{D}_{\mathrm{KL}}^{\cos}(\mu_{1},\mu_{2},\sigma_{1}^{2},\sigma_{2}^{2})$ $=\frac{1}{2}\Big{[}\log(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+(1-\cos(\mu_{1},\mu_{2}))^{2}$
$\quad+(1-\cos(\sigma_{1},\sigma_{2}))-1\Big{]}$
Expected Cosine-Based Log-Likelihood:
~$\tilde{\mathcal{L}}^{\cos}_{\text{ll}}(\mu_{1},\mu_{2},$ $\sigma_{1}^{2},\sigma_{2}^{2})=\log(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+4(1-\cos(\sigma_{1},\sigma_{2}))$
$\quad+(1-\cos(\mu_{1},\mu_{2}))^{2}(1-\cos(\sigma_{1},\sigma_{2}))$
These formulations leverage:
  • Log-term: captures similarity in uncertainty.
  • Variance contrast: emphasizes alignment of distributional spread.
  • Mean similarity: enforces semantic closeness, weighted by variance agreement.
Cosine similarity provides a robust angular measure, particularly effective in high-dimensional embedding spaces, offering an alternative to Euclidean-based alignment in variational settings.
Algorithm 1 Variational Self-Supervised Learning
0:  Dataset $\mathcal{D}=\{x^{(i)}\}_{i=1}^{N}$
0:  Trained student and teacher parameters $\phi,\theta$
1:  Initialize student $\phi$ and teacher $\theta$ with momentum copies
2:  for each epoch do
3:     for each mini-batch $\{x_{1},x_{2},\dots,x_{k}\}$ from $\mathcal{D}$ do
4:        for each view $x_{i}$ do
5:           Encode using student: $\mu_{i},\sigma_{i}=\text{Pj}_{\phi}(f(x_{i}))$
6:           Predict posterior: $\hat{\mu}_{i},\hat{\sigma}_{i}=\text{Pd}_{\phi}(\mu_{i},\sigma_{i})$
7:        end for
8:        for each view $x_{j}$ do
9:           Encode using teacher:
10:           $\mu_{j}^{\text{mom}},\sigma_{j}^{\text{mom}}=\text{Pj}_{\theta}(f(x_{j}))$
11:           Predict momentum prior:
12:           $\hat{\mu}_{j}^{\text{mom}},\hat{\sigma}_{j}^{\text{mom}}=\text{Pd}_{\theta}(\mu_{j}^{\text{mom}},\sigma_{j}^{\text{mom}})$
13:        end for
14:        Compute KL divergence between posteriors and priors: $\mathcal{L}_{\text{KL}}$
15:        Denoise student samples to get $(\mu^{\text{recon}},\sigma^{\text{recon}})$
16:        Compute expected log-likelihood loss: $\mathcal{L}_{\text{ll}}$
17:        Compute total loss: $\mathcal{L}=\mathcal{L}_{\text{KL}}+\mathcal{L}_{\text{ll}}$
18:        Update $\phi$ using gradients from $\nabla_{\phi}\mathcal{L}$
19:        Update $\theta$ using EMA of $\phi$
20:     end for
21:  end for
IV Experimental Evaluation
We evaluate the proposed VSSL method using the solo-learn framework [15] across three benchmark datasets: CIFAR-10, CIFAR-100 [16], and ImageNet-100 [17]. Model performance is assessed using both online and offline (only for ImageNet-100) linear evaluation protocols, reporting top-1 classification accuracy.
To ensure a fair and meaningful comparison, we conduct extensive hyperparameter tuning for VSSL. For the comparison table we use high-quality open-source baselines results by solo-learn. In many cases, our reproduced baselines by solo-learn achieve results that are stronger than those originally reported, enabling a more robust evaluation setting.
The full results are summarized in Table I, where we highlight the best, second-best, and third-best performances across all datasets and settings.
Method CIFAR-10 Acc@1 CIFAR-100 Acc@1 ImageNet-100 (Online) ImageNet-100 (Offline)
Barlow Twins [11] 92.10 70.90 80.38 80.16
BYOL [3] 92.58 70.46 80.16 80.32
DeepCluster V2 [18] 88.85 63.61 75.36 75.40
DINO [19] 89.52 66.76 74.84 74.92
MoCo V2+ [20] 92.94 69.89 78.20 79.28
MoCo V3 [21] 93.10 68.83 80.36 80.36
NNCLR [22] 91.88 69.62 79.80 80.16
ReSSL [23] 90.63 65.92 76.92 78.48
SimCLR [2] 90.74 65.78 77.04 -
SimSiam [8] 90.51 66.04 74.54 78.72
SwAV [24] 89.17 64.88 74.04 74.28
VICReg [12] 92.07 68.54 79.22 79.40
VSSL (this work) 93.08 70.53 81.50 81.34
TABLE I: Top-1 classification accuracy (%) of various self-supervised learning methods on CIFAR-10, CIFAR-100, and ImageNet-100 under both online and offline linear evaluation protocols. The best results are shown in bold, while the second- and third-best results are underlined. Our proposed method, VSSL, achieves competitive or state-of-the-art performance across multiple datasets and settings.
V Results and Discussion
Table I presents the top-1 classification accuracy (%) of several prominent self-supervised learning (SSL) methods across three benchmark datasets: CIFAR-10, CIFAR-100, and ImageNet-100. We evaluate each method under both online and offline linear evaluation protocols.
Our proposed method, VSSL, consistently demonstrates strong performance across all benchmarks. On CIFAR-10, VSSL achieves an accuracy of 93.08%, ranking second only to MoCo V3 (93.10%) by a narrow margin of 0.02%. On CIFAR-100, VSSL achieves 70.53%, placing it among the top three performers, slightly behind Barlow Twins (70.90%) and BYOL (70.46%).
On the more challenging ImageNet-100 dataset, VSSL outperforms all competing methods under both evaluation protocols. Specifically, it achieves 81.50% in the online setting and 81.34% in the offline setting, surpassing the closest competitors, BYOL and MoCo V3, by over 1%.
These results highlight the effectiveness of our approach in capturing transferable visual representations. VSSL not only maintains competitive performance on relatively simpler datasets (e.g., CIFAR-10/100) but also excels on large-scale, diverse datasets such as ImageNet-100, where robust representation learning is crucial.
Interestingly, methods like Barlow Twins and BYOL show strong results on CIFAR datasets but exhibit diminished performance on ImageNet-100. In contrast, approaches such as MoCo V3 and NNCLR achieve more balanced results across different data scales. VSSL appears to combine the strengths of both, suggesting better generalization and robustness.
Overall, the consistent superiority of VSSL across different datasets and evaluation settings underscores its potential as a versatile and scalable SSL framework. Further ablation studies and transfer learning experiments may provide additional insight into the key components contributing to its performance.
VI Conclusion
We presented VSSL, a novel variational self-supervised learning framework that eliminates the need for a decoder by adopting a symmetric encoder architecture and aligning representations through Gaussian latent distributions. By integrating variational inference with momentum-based contrastive learning, VSSL replaces traditional reconstruction-based objectives with a denoising cross-predictive task, maintaining analytical tractability while enabling efficient and scalable representation learning.
Empirical results across CIFAR-10, CIFAR-100, and ImageNet-100 demonstrate that VSSL matches or surpasses state-of-the-art self-supervised learning methods, highlighting its robustness and generalization capabilities. Beyond performance, VSSL contributes a new perspective to the design of latent variable models by bridging probabilistic modeling and self-supervised learning—without relying on generative reconstruction. This work opens new directions for lightweight, probabilistically grounded SSL methods that are both theoretically sound and practically effective.
Acknowledgment
The numerical calculations reported in this paper were fully/partially performed at TUBITAK ULAKBIM, High Performance and Grid Computing Center (TRUBA resources).